OpenStudy (anonymous):

(-2,-3) & (-6,-5) answer is in slope-intercept form please and thank you in advance

OpenStudy (anonymous):

let (-2, -3) = (x1, y1) and let (-6, -5) = (x2,y2) slope = m = (y2-y1)/(x2-x1) = (1/2) point slope form: y-y1 = m(x-x1) y-(-3) = (1/2)(x-(-2)) y+3 = (1/2)x + 1 slope int form = y = (1/2)x -2

OpenStudy (anonymous):

you want it in the form \[y = mx + b\] so you do this by first finding m the equation is m = (y2-y1)/(x2-x1) so just use the points you are given m=(-5-(-3))/(-6-(-2)) = (-2)/(-4) = 1/2 so now solve for b using the equation and plugging in m and x and y so use (-2, -3) -3 = (1/2)(-2) + b -3 = -1 + b -2 = b then now plug in m and b for the general equation y = (1/2)x - 2